When I tried to solve the sample problem my answer was wrong 2x-1y=9 5x+5y=0 this is what I did 2x-1y=9
5x+5y=0
7x+4y=9 for substitution?

Answer by Alan3354(69443)

When I tried to solve the sample problem my answer was wrong 2x-1y=9 5x+5y=0 this is what I did 2x-1y=9
5x+5y=0
7x+4y=9 for substitution?
-----------
Using substition:
2x-1y=9
5x+5y=0 --> y = -x
Sub for y in the 1st eqn
2x - (-x) = 9
3x = 9
x = 3
y = -3
==========
Using elimination:
2x-1y=9
5x+5y=0 --> x + y = 0
------
2x- y = 9
x + y = 0
------------- Add
3x = 9
etc.